# Domain-Invariant Representation Learning of Bird Sounds

### v5

@@ IV-D Results and Discussion
For self-supervised methods, SimCLR outperforms BirdAVES-biox-base and BirdAVES-bioxn-large in both one-shot and five-shot learning, suggesting that invariant learning is more effective than self-prediction methods for few-shot scenarios, consistent with findings in computer vision [35].
@@
On average, BioLingual outperforms CE in one-shot learning but underperforms in five-shot learning. However, when examining individual datasets, BioLingual performs worse on PER, NES, UHH, and HSN, yet significantly better on SSW and SNE. Interestingly, certain models appear to specialize in specific species, suggesting a promising research direction: distilling these specialized models into a single general model that can leverage the strengths of each.
@@
Perch significantly outperforms all models in both one-shot and five-shot classification across all datasets, except for PER, where it performs slightly worse than CE. This suggests that incorporating taxonomic ranks as auxiliary tasks may enhance the discriminative capabilities of the feature extractor.
V Conclusion
@@
In this work, we addressed the challenge of domain generalization for bird sound classification in few-shot scenarios, with a focus on the domain shift from focal to soundscape recordings. We proposed a new few-shot evaluation, derived from the BIRB datasets, to evaluate the generalization capabilities of models trained on focal recordings and tested on soundscapes. Additionally, we introduced ProtoCLR, a computationally efficient alternative to SupCon, inspired by prototypical networks. ProtoCLR maintains performance levels comparable to SupCon, and in some cases, slightly outperforms it.
Future work will explore scaling training to longer epochs to maximize the potential of contrastive methods, as well as integrating taxonomic ranks into the learning process. Another promising direction is to investigate domain adaptation techniques, particularly source-free adaptation, which is well-suited for few-shot bioacoustics when access to source data is limited or retraining is too costly.

### v6

@@ 4.4 Results and Discussion
For self-supervised methods, SimCLR outperforms BirdAVES-biox-base and BirdAVES-bioxn-large in both 1-shot and 5-shot learning, despite using a smaller model and simple data augmentation. This suggests that invariant learning is more effective than self-prediction methods for few-shot scenarios, consistent with findings in computer vision [37].
@@
On average, ProtoCLR outperforms BioLingual in both 1-shot learning and 5-shot learning. However, when examining individual datasets, BioLingual performs worse on PER, NES, UHH, and HSN, yet significantly better on SSW and SNE. Interestingly, certain models appear to specialize in specific species, suggesting a promising research direction: distilling these specialized models into a single general model that can leverage the strengths of each.
@@
Perch outperforms all models in both 1-shot and 5-shot classification across all datasets, except for PER, where it performs slightly worse than CE. This suggests that incorporating taxonomic ranks as auxiliary tasks can enhance the discriminative capabilities of the feature extractor.
5 Conclusion
@@
In this work, we addressed the challenge of domain generalization for bird sound classification in few-shot scenarios, with a focus on the domain shift from focal to soundscape recordings. We proposed a new few-shot evaluation, derived from the BIRB datasets, to evaluate the generalization capabilities of models trained on focal recordings and tested on soundscapes. Additionally, we introduced ProtoCLR, a computationally efficient alternative to SupCon, inspired by prototypical learning. ProtoCLR maintains performance levels comparable to SupCon, and in some cases, outperforms it.
SimCLR is shown to outperform state-of-the-art self-supervised models. It falls behind supervised models but offer a scaling advantage when large unlabeled data is available. Without strong data augmentation, CE outperforms supervised contrastive methods. Future work will explore data augmentation for bioacoustics to efficiently leverage invariant learning. Incorporating taxonomic ranks into the learning process holds promise as shown by Perch. One potential limitation of ProtoCLR, as well as other state-of-the-art approaches, is the assumption that each example belongs to a single class. This is often not the case in bioacoustics, where sounds can overlap in soundscape recordings. Future work should consider an extension of ProtoCLR that allows examples to belong to multiple classes.